OpenStudy (anonymous):

Find the least value of the expression E=3x^2+5y^2-6x-30y+50

OpenStudy (anonymous):

@mukushla @RadEn @zepdrix

OpenStudy (anonymous):

make 2 complete squares...1 in terms of x another in terms of y

OpenStudy (anonymous):

E=3x^2+5y^2-6x-30y+50 3 (x^2 - 2x + 1) + 5 ( y^2-6y + 9) + 50 - 3 -45

OpenStudy (anonymous):

3 (x-1)^2 + 5(y-3)^2 + 2 =E

OpenStudy (anonymous):

Then ......?

OpenStudy (anonymous):

well minimum is 2 :) right??

OpenStudy (anonymous):

Yup./..u r ryt ? but how ?

OpenStudy (raden):

yeah, minimum at poit (1,3), it is 2

OpenStudy (anonymous):

\[ (x-1)^2 \ge0\]\[(y-3)^2 \ge 0\]so\[3 (x-1)^2 + 5(y-3)^2 + 2 \ge 0+0+2=2\]notice that 2 squares are independent of each other

OpenStudy (anonymous):

Errr.......Thxxx//..

OpenStudy (anonymous):

welcome :) tnx raden for medal man :)

OpenStudy (anonymous):

Nice solution!